# definitionDefinition[section]

### v1

@@ 1 Introduction
The use of high-order discontinuous Galerkin (DG) schemes for simulating transport-dominated physics often requires additional numerical stabilization to ensure the discrete solution is structure-preserving (i.e., is well-behaved and satisfies physical constraints). One such way of achieving this desirable behavior is applying some form of a posteriori limiting on the solution, where the high-order DG approximation is blended with a secondary, more robust (but generally less accurate) approximation to yield a bounds-preserving (i.e., constraint-satisfying) solution. This form of limiting has been widely used for simulating complex physical systems, particularly ones of predominantly hyperbolic nature, where numerical solutions must abide by known physical constraints (e.g., positivity of density and pressure in gas dynamics). However, a potential drawback of these limiting methods is that they are typically performed at discrete nodal locations for the DG solution, which only ensures that the given constraints are satisfied at those points. While adequate for many problems, the use of discretely bounds-preserving limiting may not be robust enough for applications where the DG solution must be evaluated at arbitrary locations (e.g., coupled meshes/solvers, remapping in arbitrary Lagrangian-Eulerian methods, adaptive mesh refinement, etc.). In such scenarios, the limited solution may still violate the constraints at these arbitrary points, causing the failure of the numerical scheme.
@@
A potential remedy for this problem was introduced in Dzanic [1] through a limiting approach that ensures the limited solution is continuously bounds-preserving (i.e., across the entire solution polynomial). The approach relied on a novel functional for computing the limiting factor for the “squeeze” limiter of Zhang and Shu [2], such that if the minimum of this limiting functional within an element was found (via a spatial optimization algorithm), the limited solution was guaranteed to be continuously bounds-preserving for any arbitrary quasiconcave constraint functional. The underlying mathematical basis of this approach was a linearization of the constraint functionals being enforced on the solution. For linear constraints (e.g., maximum principle on scalar solutions, positivity of density in gas dynamics, etc.), the computed limiting factor was “exact” (i.e., it was the minimum necessary amount of limiting). However, for nonlinear constraint functionals, the linearization could only ensure that the computed limiting factor was sufficient, such that a smaller limiting factor (i.e., less numerical dissipation) could still yield a solution which satisfied the constraints continuously. Therefore, the limiting approach was suboptimal for nonlinear constraints commonly encountered in hyperbolic systems of equations as it was not applying only the minimum amount of limiting necessary.
@@ 2.1 Governing equations and constraints
Here, $\rho$ is the density, $\mathbf{m}$ is the momentum, and $E$ is the total energy. The symbol $\mathbf{I}$ denotes the identity matrix in $\mathbb{R}^{d\times d}$, $\mathbf{v}=\mathbf{m}/\rho$ denotes the velocity, and $P$ denotes the pressure, computed as
@@ 2.2 Discontinuous Galerkin methods
The underlying numerical method for this approach is the discontinuous Galerkin scheme [4], where the solution $\mathbf{u}(\mathbf{x})$ within each element of the mesh is represented by a set of $n_{s}$ basis functions as
@@
$\overline{\mathbf{u}}=\frac{\int_{\Omega_{k}}\mathbf{u}(\mathbf{x})\ \mathrm{d}\mathbf{x}}{\int_{\Omega_{k}}\mathrm{d}\mathbf{x}},$ (7)
@@ 2.3 Continuously bounds-preserving limiting
The goal of the proposed approach is to find a limiting factor such that
@@ 3 Proposed modifications
To this end, we propose achieving this by replacing the linearized formulation of $h^{-}(\mathbf{u})$ is Eq. 10, denoted by $h^{-}_{L}(\mathbf{u})$, with a better (nonlinear) approximation of the necessary limiting factor, denoted by $h^{-}_{NL}(\mathbf{u})$, which mimics the above condition as
@@
for positive pressure constraints, where $\Delta\mathbf{u}=\overline{\mathbf{u}}-\mathbf{u}$. Similar closed-form expressions can be used for higher-order constraints (e.g., cubic) which admit analytic solutions to the intersection problem in Eq. 13.
@@ 4 Results
The modified functional was implemented within the same numerical framework as in Dzanic [1], using identical parameters and optimization approaches (for more information, the reader is referred to Section 3 of Dzanic [1]). The intersection problem for entropy constraints was solved using 5 iterations of the Illinois method. Minimum density and pressure values were set as $\rho_{\min}=P_{\min}=10^{-11}$.
@@
A 9th-order polynomial DG approximation was initialized by interpolating the solution on the Gauss–Lobatto nodes spaced along the domain $\Omega=[0,1]$, and constraints were enforced for positivity of density/pressure as well as a minimum entropy principle. An arbitrary minimum entropy value was set as $\sigma_{\min}=0.1$. These conditions yield various properties for the unlimited solution to showcase the behavior of the new approach, with a continuously bounds-preserving density field, a discretely (but not continuously) bounds-preserving pressure field, and a discretely bounds-violating entropy field.
@@
The parameters are set to $R=1.5$, $M=0.4$, and $S=28.11711$, which yield a near-vacuum state at the vortex peak with a minimum density of $\rho=8{\cdot}10^{-9}$ and minimum pressure of $P=2{\cdot}10^{-11}=2P_{\min}$. One flow-through of the domain ($t=20$) was computed using both a $\mathbb{P}_{4}$ and $\mathbb{P}_{5}$ approximation using meshes of varying resolution with only positivity-preserving constraints, after which the $L^{\infty}$ norm of the pressure error (computed at solution nodes) was compared from the solutions obtained using the original linear limiting approach and the proposed nonlinear limiting approach. The comparisons for both the $\mathbb{P}_{4}$ and $\mathbb{P}_{5}$ approximations are shown in Section 4 and Section 4, respectively. It can be seen that the proposed modification significantly reduces the overall error in the pressure field, with both approximation orders showing decreases between $20-70\%$ across the varying levels of mesh resolution. These results highlight the reduced numerical dissipation that stems from the modified constraint functional proposed in this note.
@@ Acknowledgements
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract DE–AC52–07NA27344 and the LLNL-LDRD Program under Project tracking No. 24–ERD–050. Release number LLNL–JRNL–863292–DRAFT.

### v2

@@ 1 Introduction
The use of high-order discontinuous Galerkin (DG) schemes for simulating transport-dominated physics often requires additional numerical stabilization to ensure the discrete solution is structure-preserving (i.e., is well-behaved and satisfies physical constraints). One such way of achieving this desirable behavior is applying some form of a posteriori limiting on the solution, where the high-order DG approximation is blended with a secondary, more robust (but generally less accurate) approximation to yield a bounds-preserving (i.e., constraint-satisfying) solution, a few examples of which are presented in the following references [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. This form of limiting has been widely used for simulating complex physical systems, particularly ones of predominantly hyperbolic nature, where numerical solutions must abide by known physical constraints (e.g., positivity of density and pressure in gas dynamics). However, a potential drawback of these limiting methods is that they are typically performed at discrete nodal locations for the DG solution, which only ensures that the given constraints are satisfied at those points. While adequate for many problems, the use of discretely bounds-preserving limiting may not be robust enough for applications where the DG solution must be evaluated at arbitrary locations (e.g., coupled meshes/solvers, remapping in arbitrary Lagrangian-Eulerian methods, adaptive mesh refinement, etc.). In such scenarios, the limited solution may still violate the constraints at these arbitrary points, causing the failure of the numerical scheme.
@@
A potential remedy for this problem was introduced in Dzanic [12] through a limiting approach that ensures the limited solution is continuously bounds-preserving (i.e., across the entire solution polynomial). The approach relied on a novel functional for computing the limiting factor for the “squeeze” limiter of Zhang and Shu [1], such that if the minimum of this limiting functional within an element was found (via a spatial optimization algorithm), the limited solution was guaranteed to be continuously bounds-preserving for any arbitrary quasiconcave constraint functional. This approach differed from other continuously bounds-preserving methods which rely on bounded basis functions (e.g., Bernstein polynomials) [3, 13, 5], which may suffer from numerical inefficiencies and excessive dissipation (a more in-depth discussion on these issues is presented in Dzanic [12], Section 1). The underlying mathematical basis of this approach was a linearization of the constraint functionals being enforced on the solution. For linear constraints (e.g., maximum principle on scalar solutions, positivity of density in gas dynamics, etc.), the computed limiting factor was “exact” (i.e., it was the minimum necessary amount of limiting). However, for nonlinear constraint functionals, the linearization could only ensure that the computed limiting factor was sufficient, such that a smaller limiting factor (i.e., less numerical dissipation) could still yield a solution which satisfied the constraints continuously. Therefore, the limiting approach was suboptimal for nonlinear constraints commonly encountered in hyperbolic systems of equations as it was not applying only the minimum amount of limiting necessary.
@@ 2.1 Governing equations and constraints
Here, $\rho$ is the density, $\mathbf{m}$ is the momentum, and $E$ is the total energy. The symbol $\mathbf{I}$ denotes the identity matrix in $\mathbb{R}^{d\times d}$, $\mathbf{v}=\mathbf{m}/\rho$ denotes the velocity, and $P$ denotes the pressure, computed (with the assumption of a calorically perfect gas) as
@@ 2.2 Discontinuous Galerkin methods
The underlying numerical method for this approach is the discontinuous Galerkin scheme (see, for example, Hesthaven and Warburton [15]), where the solution $\mathbf{u}(\mathbf{x})$ within each element of the mesh is represented by a set of $n_{s}$ polynomial basis functions as
@@
$\overline{\mathbf{u}}_{k}=\frac{\int_{\Omega_{k}}\mathbf{u}(\mathbf{x})\ \mathrm{d}\mathbf{x}}{\int_{\Omega_{k}}\mathrm{d}\mathbf{x}},$ (7)
@@ 2.3 Continuously bounds-preserving limiting
We drop the subscript $k$ for brevity as this limiter applies to every element in the domain. The goal of the proposed approach is to find a limiting factor such that
@@ 3 Proposed modifications
To achieve this, we propose replacing the linearized formulation of $h^{-}(\mathbf{u})$ in Eq. 10, denoted by $h^{-}_{L}(\mathbf{u})$, with a better (nonlinear) approximation of the necessary limiting factor, denoted by $h^{-}_{NL}(\mathbf{u})$, which mimics the above condition as
@@
for positive pressure constraints, where $\Delta\mathbf{u}=\overline{\mathbf{u}}-\mathbf{u}$. This quadratic formulation is similarly seen in the works of Kuzmin and Lohmann [2] and Zhang and Shu [16]. Similar closed-form expressions can be used for higher-order constraints (e.g., cubic) which admit analytic solutions to the intersection problem in Eq. 13.
@@
4 Implementation
The modified functional was implemented within the same numerical framework as in Dzanic [12] and closely follows the original work, using identical parameters and optimization approaches. This approach is summarized as follows, but for more information, the reader is referred to Section 3 of Dzanic [12]. The solution was first converted into a modal basis in monomial form for fast evaluation at arbitrary locations. The initial guess for constraint functional for the optimization process was evaluated at the solution nodes (e.g., Gauss–Lobatto nodes) as well as the volume/surface quadrature nodes (e.g., Gauss–Legendre nodes), the minimum of which was taken as the initial point for optimization. We note that the distinction between solution nodes and quadrature nodes is not necessarily present in collocated flux reconstruction schemes with closed solution nodes, but we evaluate at these nodes regardless to mimic standard nodal DG approaches. Optimization was performed using two iterations of the Newton–Raphson method with a fallback on adaptive gradient descent with backtracking line search in regions where Newton–Raphson is ill-suited. The Jacobian and Hessian of the constraint functionals were computed numerically, and a bound on the true minimum of the constraint functional was extrapolated at the end of the optimization steps. For the gas dynamics equations, the constraints were enforced sequentially, first on density, then pressure, then entropy, to ensure that the constraint functionals were well-behaved. Minimum density and pressure values were set as $\rho_{\min}=P_{\min}=10^{-11}$. A minimum tolerance was set as $\epsilon=10^{-12}$, and the limiting factor was explicitly set as $\alpha=1$ when $g(\overline{\mathbf{u}})<\epsilon$ to avoid numerically undefined behavior.
Some further implementation details are necessary for the proposed modification to the approach. For pressure constraints, it is necessary to compute $h_{NL}^{-}(\mathbf{u})$ as the root of the quadratic polynomial. It can be shown that in locations where $h(\mathbf{u})=h_{NL}^{-}(\mathbf{u})$ (i.e., where $g(\mathbf{u})<0$) and under the assumption that $g(\overline{\mathbf{u}})>0$, the discriminant of the quadratic equation is positive and only one of the roots, given by
$\frac{-B+\sqrt{B^{2}-4AC}}{2A},$
yields a limiting factor $\alpha^{*}\in[0,1]$. In practice, care must be taken to avoid numerical round-off errors in calculating this root, namely by enforcing a floor value of zero for the discriminant. The edge case of $A\sim 0$ is taken care of in a similar manner to the edge case of $g(\overline{\mathbf{u}})\sim 0$, where the limiting factor is explicitly set as $\alpha^{*}=1$ when $|A|<\epsilon$. Therefore, for computational purposes, the limiting factor is computed as
$\alpha^{*}=\begin{cases}1,\hskip 82.0pt\text{if }|A|<\epsilon,\\ \frac{-B+\sqrt{\max(0,B^{2}-4AC)}}{2A},\quad\text{else}.\end{cases}$ (14)
For the entropy constraints, 5 iterations of the Illinois method were used to compute $h_{NL}^{-}(\mathbf{u})$ in locations where $g(\mathbf{u})<0$. The initial lower (bounds-violating) bracket was set as $\alpha_{l}=0$ with $g_{l}=g(\alpha_{l})$, and the upper (bounds-preserving) bracket was set as $\alpha_{h}=-h_{L}^{-}(\mathbf{u})$ with $g_{h}=g(\alpha_{h})$ (recall that the linearized formulation always yields an upper bound on the exact limiting factor necessary). At each iteration, the next guess was computed as
$\alpha_{m}=\frac{\alpha_{l}g_{h}-\alpha_{h}g_{l}}{g_{h}-g_{l}},$ (15)
and the brackets were updated as
$\alpha_{l}=\alpha_{m},\quad g_{l}=g(\alpha_{m}),\quad\text{and}\quad g_{h}=\frac{1}{2}g(\alpha_{h})$
if $g(\alpha_{m})<0$ and
$\alpha_{h}=\alpha_{m},\quad g_{l}=\frac{1}{2}g(\alpha_{l}),\quad\text{and}\quad g_{h}=g(\alpha_{m})$
if $g(\alpha_{m})\geq 0$. At the end of the root-finding process, the solution was set to the upper bound, i.e., $\alpha^{*}=-\alpha_{h}$, to ensure that the bounds-preserving estimate of the limiting factor was chosen.
@@ 5 Results
A 9th-order polynomial DG approximation was initialized by interpolating the solution on the Gauss–Lobatto nodes spaced along the domain $\Omega=[0,1]$, and constraints were enforced for positivity of density/pressure as well as a minimum entropy principle. An arbitrary minimum entropy value was set as $\sigma_{\min}=0.1$. These conditions yield various properties for the unlimited solution to showcase the behavior of the new approach, with a continuously bounds-preserving density field, a discretely (but not continuously) bounds-preserving pressure field, and a discretely bounds-violating entropy field. The minimum entropy value was chosen to demonstrate the nonlinear limiting approach for bounds-violating solution as well as to mimic a condition encountered in time-dependent simulations of hyperbolic conservation laws, where entropy bounds (which might be computed from the previous time step) are violated by the discrete solution at the next time step [6, 11].
@@
The parameters are set to $R=1.5$, $M=0.4$, and $S=28.11711$, which yield a near-vacuum state at the vortex peak with a minimum density of $\rho=8{\cdot}10^{-9}$ and minimum pressure of $P=2{\cdot}10^{-11}=2P_{\min}$. One flow-through of the domain ($t=20$) was computed using both a $\mathbb{P}_{4}$ and $\mathbb{P}_{5}$ approximation using meshes of varying resolution with only positivity-preserving constraints, after which the $L^{\infty}$ norm of the pressure error (computed at solution nodes) was compared from the solutions obtained using the original linear limiting approach and the proposed nonlinear limiting approach. The comparisons for both the $\mathbb{P}_{4}$ and $\mathbb{P}_{5}$ approximations are shown in Section 5 and Section 5, respectively. It can be seen that the proposed modification significantly reduces the overall error in the pressure field, with both approximation orders showing decreases between $20-70\%$ across the varying levels of mesh resolution. These results highlight the reduced numerical dissipation that stems from the modified constraint functional proposed in this note. Furthermore, the imposition of exact pressure constraints (as opposed to a linearized approximation) can have further benefits in terms of the computational cost as the reduced pressure (i.e., reduced speed of sound) can increase the maximum admissible explicit time step for near-vacuum problems. A simple computational cost comparison (at a fixed time step) was performed for the near-vacuum isentropic Euler vortex problem using a $\mathbb{P}_{4}$ approximation with $N=30^{2}$ elements on one NVIDIA V100 GPU in terms of the absolute wall clock time. As the added effort of analytically computing $h_{NL}^{-}(\mathbf{u})$ was essentially negligible compared to computing $h_{L}^{-}(\mathbf{u})$, the computational cost of the approach with the proposed modification was effectively identical to the linearized method, both with a computational cost increase of approximately $18\%$ over a discrete limiting approach. When $h_{NL}^{-}(\mathbf{u})$ was instead computed numerically (which is not necessary for the pressure constraints used for this example but is simply presented for comparison), the overall computational cost increase in comparison to a discrete limiting method was then $24.7\%$. We remark here that the relatively marginal difference between the analytic and numerical nonlinear approaches can largely be attributed to the efficacy of GPU computing for local compute-intensive operations.
@@ Acknowledgements
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract DE–AC52–07NA27344 and the LLNL-LDRD Program under Project tracking No. 24–ERD–050. Release number LLNL–JRNL–863292.